Using 20 observations, the following regression output is obtained from estimating y = β0 + β1x + β2d + β3xd + ε.

Coefficients Standard Error t Stat p-value
Intercept 13.05 3.00 4.35 0.001
x 3.76 0.47 8.00 0.000
d −4.59 3.06 −1.50 0.153
xd 1.89 0.70 2.70 0.016


a. Compute yˆy^ for x = 11 and d = 1; then compute yˆy^ for x = 11 and d = 0. (Round intermediate calculations to at least 4 decimal places and final answers to 2 decimal places.)




b-1. Is the dummy variable d significant at the 5% level?

  • Yes, since we reject the relevant null hypothesis.

  • Yes, since we do not reject the relevant null hypothesis.

  • No, since we reject the relevant null hypothesis.

  • No, since we do not reject the relevant null hypothesis.



b-2. Is the interaction variable xd significant at the 5% level?

  • Yes, since we reject the relevant null hypothesis.

  • Yes, since we do not reject the relevant null hypothesis.

  • No, since we reject the relevant null hypothesis.

  • No, since we do not reject the relevant null hypothesis.
Answer #1

a)

for x=11 and d=1 ;y^ =13.05+3.76*11-4.59*1+1.89*11*1 =70.61

for x=11 and d=0 ;y^ =13.05+3.76*11-4.59*0+1.89*11*0 =54.41

b-1)since p value >0.05

  • No, since we do not reject the relevant null hypothesis.

b-2)soince for interaction : p value 0.016 < 0.05

es, since we reject the relevant null hypothesis.